# Fast Relax-and-Round Unit Commitment with Sub-hourly Mechanical and Ramp Constraints

### 04 Mar 2026 v2

@@ I Introduction
Generator EIA data shows they are becoming smaller. For units in operation in 2024, excluding Alaska and Hawaii, average nameplate capacity for a generator built in 1974 was 205.22 MW. For a generator built in 2024 it was 49.5 MW [8]. For 2408 generating units listed in “proposed”, the average nameplate capacity was 80.21 MW for units dated 2014-2024 and 193.83 MW for units planned for 2025-2030. However, the actual average capacity is expected to be lower. First, historically, proposed units between 2014 and 2024 were approximately 1.5-4.5 times larger than those ultimately built.
@@ II-A Mathematical Formulation
  • $\sigma_{D}$: standard deviation (STD) of demand forecast
@@ III-A Mathematical Formulation
$\min\quad$ $\sum_{i\in G_{m}}\left(a_{i}P_{i}^{2}+b_{i}P_{i}+c_{i}\right)$
$+\sum_{i\in G_{d}}\left[\left(a_{i}P_{i}^{2}+b_{i}P_{i}+c_{i}\right)u_{i}+K_{i}(u_{i}-u_{i}^{0})^{2}\right]$ (2a)
@@
Note that at the lower end, we plan for a smaller contingency ($1\sigma_{D}$) than on the higher end ($3\sigma_{D}$). This is meant to reflect the asymmetric nature of load forecasts. It is far more likely that load will be much higher than predicted (corresponding to some large unexpected load), compared to much lower (corresponding to base load being removed). This also helps provide additional buffers in the event that more than 1 generator goes out, or uncontrollable generation is much lower than expected.
@@
  • $P_{\max,i},P_{\min,i}$: maximum capacity and minimum stable output, usually determined by mechanical characteristics of the units.
@@
  • $K_{i}$: commitment change penalty for generator $i$. The true costs of startup and shutdown are usually not the same. Averaging the startup and shutdown costs prevents market distortions and ensures that there are no artificial incentives to run.
  • $u_{i}^{0}\in\{0,1\}$: previous commitment status of generator $i$
@@
$\min\quad$ $\sum_{j=1}^{k}\left(a_{j}P_{j}^{2}+b_{j}P_{j}+c_{j}+K_{j}(1-u_{j}^{0})\right)$ (5a)
@@
1:for $i=1$ to number_of_periods do
2:  Update rolling 24 hour generator states (use them in steps 3 and 4).
3:  Find must run generators (those whose minimum on time has not been reached).
4:  Find can’t run generators (those who have exceeded the maximum allowed daily starts).
@@
6:  for Different sets of discretionary generators sufficient to meet the constraints (using Eq. 3 and Eq. 4) do
7:   Solve economic dispatch Eq. 5 with all must run generators and this discretionary generator set.
8:   if Objective is lowest to far then
9:    Update best solution and objective.
10:   end if
11:  end for
12:  return Best solution and best objective.
13:end for
@@
The results of this algorithm would not be directly comparable to fully coupled approach on computational time. Since we have shown our algorithm scales sub-quadratically [27], we believe it is possible to proceed with this approach and expand it in the future to include a fully time-coupled objective. In this work, the runtime constraints are still time-coupled and the system would be forced to converge while fully accounting for minimum runtime or the number of starts.
Unlike [27], we do not compare RRUC results to MIP results. The large system sizes, combined with new constraints, make MIP prohibitively computationally intensive for the available hardware (Apple M3 Pro, 18GB RAM). Instead we focus specifically on RRUC and adopt a new metric to assess the efficiency of the algorithm: percentage increase in the objective function and percentage increase in runtime as a function of increase in the system size. If the objective function and runtime do not explode as the system size increases. The algorithm maintains its scaling properties and is efficient.
@@ III-B Simulation and Results
We multiply this generator set by 22 to pass 192 GW, using a random seed to rescale each generator in order to avoid the symmetry effects [19]. Non-identical copies of the original 42 generators have a shift in capacity parameters of $\pm 10\%$ and a shift in daily starts or runtime hour parameters of $\pm 1$. For various system sizes, we further use multiples of this system.
@@
We present the results of the simulation in Fig. 2 using the same software and hardware as before. Since the reference MIP solver is too slow to be used as a benchmark for RRUC, in this study we are mainly interested in analyzing the scaling and performance of RRUC. The runtime and the objective function increase slightly faster than linearly with the number of generators. Specifically, the runtime increases on average about $2.6\times$ for every doubling of the system size. The $\approx O(n^{1.37})$ shows that the resulting systems are less complicated than those arising from a typical optimization problem which have $O(n^{1.5})$ scaling. For a general optimization problem, $O(n^{1.5})$ scaling is believed to be ideal, due to the bottleneck from the linear solver [20].
@@ IV-A Mathematical Formulation
We now extend the problem to larger and more varied thermal power systems, which have have non-thermal units with intermittent characteristics. To operate such systems, we must account for system behavior at a higher resolution, which requires a sub hour granularity of unit commitment. As discussed in Section I, high temporal resolution involves another important aspect of generator behavior, ramp.
@@
To address the ramp constraints in the model, we use a different set of variables. Instead of the original $u_{i}$ that indicated that a generator is on or off, we allow the generator to be off, preparing (for ramp up), ramping up, on, or ramping down. For the model purposes, we assume that a generator can only change its status in one direction. If a generator is ramping up or preparing, it is assigned a variable $v_{i}=1$. After the required ramp time in each stage, its status will switch to on. If a generator is ramping down, it is assigned a variable $w_{i}=1$. It will eventually switch off and enter the set of generators that can start.
@@
$+\sum_{i\in G_{d}}\left[\left(a_{i}P_{i}^{2}+b_{i}P_{i}+c_{i}\right)u_{i}+K_{i}(w_{i}+v_{i})^{2}\right]$
@@
  • $v_{i}\in\{0,1\}$ for $i\in\mathcal{G}_{d}$ when the generator was previously off. 0 represents continuing to stay off, $1$ represents starting to ramp up.
  • $w_{i}\in\{0,1\}$ for $i\in\mathcal{G}_{d}$ when the generator was previously on. 0 represents continuing to stay on, $1$ represents starting to ramp down.
  • $u_{i}\in\{0,1\}$ for $i\in\mathcal{G}_{d}$: commitment status of discretionary generator $i$. We eliminate this variable in practice because it can be deduced from the previous status, $v_{i}$ and $w_{i}$.
@@
  • $P_{\max,i},P_{\min,i}$: maximum capacity and minimum stable output, usually determined by physical characteristics of the units.
@@
  • $K_{i}$: commitment change penalty for generator $i$. The true costs of startup and shutdown are usually not the same. Averaging the startup and shutdown costs prevents market distortions and ensures that there are no artificial incentives to run.
  • $r_{d,i}$ and $r_{u,i}$, generator $i$’s limits for ramp down or up in a given time period.
  • $\beta$ parameter we set to $0.001$. The corresponding term in the objective pushes us to choose more efficient generators for meeting the contingencies.
  • $P_{typ,i}$, a typical production level of a generator during a peak event, which we define as $(4*P_{\max,i}+P_{\min,i})/5$.
@@
  • $S_{r}$ - supply produced by generators that are already ramping before the solve (they cannot be changed). For generators ramping up, they are assumed to generate nothing for some number of periods, and then increase linearly to $P_{\min,i}$ in some number of time periods specific to the generator. Generators ramping down are assumed to decrease linearly from $P_{\min,i}$ in some number of time periods specific to the generator.
  • $S_{\max,r}$ total capacity of all generators ramping up (ramping down generators should be off by the time we reach the contingency)
  • $S_{\min,r}$ minimum operational capacity of all generators ramping up
@@
3:  Increment the number of ramping periods for all ramping generators.
4:  Switch generators that have finished preparation to ramping up.
5:  Switch generators that have finished ramping up to on.
6:  Switch generators that have finished ramping down to off.
7:  Find must run generators (those whose minimum on time has not been reached or are producing too far above $P_{\min,i}$ to shut off).
8:  Find can’t start generators (those who have exceeded the maximum allowed daily starts or are already ramping).
9:  Solve Eq. 6 with $u_{i}\in\{0,1\}$ relaxed to $0\leq y_{i}\leq 1$.
10:  for Different sets of discretionary generators sufficient to meet the constraints (using Eq. 3 and Eq. 4) do
11:   Solve economic dispatch Eq. 5 with all must run generators and this discretionary generator set.
12:   if Objective is lowest so far then
13:    Update best solution and objective.
14:   end if
15:  end for
16:  return Best solution and best objective.
17:end for
@@ IV-B Simulation and Results
Much of the data used in this Section has already been presented in Section III for runtime constraints. The additional assumptions about equipment characteristics are presented in Table II. While ramp characteristics usually come in MW/min, we use the % of $P_{\max,i}$ per minute as the original input into the simulation. The percentage values are compiled from several sources, including the academic reports from Deutsches Institut für Wirtschaftsforschung [3] and manufacturer catalog of GE Vernova [13]. The provided numbers, although summarising different sources, may still present the upper bound of ramp rates, as comprehensive survey of large thermal units suggest that in real systems power plants do not always deliver the nominal ramp rates specified for equipment [25].
@@
Fig. 4 summarizes the runtime and objective results. These results are comparable to those found in Section III using the same software and hardware. The runtime scaled roughly as $n^{1.6}$, where $n$ is the number of generators. Ideal scaling is believed to be bounded below by $n^{1.5}$, due to the scaling of the underlying linear system solvers [20]. This suggests that our optimization solver has close to ideal scaling. It is well below the exponential cost of MIP solvers. The average cost of the system did not increase with the system size. So adding extra generators does not result in scale cost effects.
@@ V Unit Commitment With Smooth Ramp Up
Figure 5: A typical production cycle for generator $i$ in Section V. When the generator is on, it can change its production power between $P_{\min,i}$ and $P_{\max,i}$, provided it does not exceed its ramping power. It can also turn off if it is close enough to $P_{\min,i}$ compared to its ability to ramp down. The only other decision can be made when the generator is off, start ramping up. During all other periods, the generator is uncontrollable.
@@ VI Conclusion and Further Work
We proposed an algorithm to solve unit commitment with mechanical and ramping constraints by relaxing the integer constraints and then rounding solutions to satisfy the constraints. Our algorithm shows between $\approx O(n^{1.6})$ time scaling in the number of generators. As the problem size increases, it retains accuracy nearly perfectly.

### 13 Mar 2026 v3

@@ I Introduction
Generator EIA data shows they are becoming smaller. For US units in operation in 2024, excluding Alaska and Hawaii, average nameplate capacity for a generator built in 1974 was 205.22 MW. For a generator built in 2024 it was 49.5 MW [8]. For 2408 generating units listed in “proposed”, the average nameplate capacity was 80.21 MW for units dated 2014-2024 and 193.83 MW for units planned for 2025-2030. However, the actual average capacity is expected to be lower. First, historically, proposed units between 2014 and 2024 were approximately 1.5-4.5 times larger than those ultimately built.
@@ II-A Mathematical Formulation
  • $\sigma_{D}$: standard deviation of demand forecast
@@ III-A Mathematical Formulation
$\min\quad$ $\sum_{i\in G_{m}}\left(a_{i}P_{i}^{2}+b_{i}P_{i}+c_{i}\right)+$ (2a)
$\sum_{i\in G_{d}}\left[\left(a_{i}P_{i}^{2}+b_{i}P_{i}+c_{i}\right)u_{i}+K_{i}(u_{i}-u_{t-1,i})^{2}\right]$
@@
  • $P_{\max,i},P_{\min,i}$: maximum capacity and minimum stable output, determined by a unit’s physical characteristics
@@
  • $K_{i}$: commitment change penalty for generator ii [^2] The true startup and shutdown costs are rarely the same. Averaging them prevents market distortion and removes artificial incentives to turn on or off..
  • $u_{t-1,i}\in\{0,1\}$: previous commitment status of unit $i$
@@
$\min\quad$ $\sum_{j=1}^{k}\left(a_{j}P_{j}^{2}+b_{j}P_{j}+c_{j}+K_{j}(u_{j}-u_{t-1,j})^{2}\right)$ (5a)
@@
1:for $i=1$ to $\mathrm{number\_of\_periods}$ do
2:  Update rolling 24-hour generator states (use them in steps 3 and 4).
3:  Find must run units (with minimum on time not yet reached or with $P_{t-1,i}-P_{min,i}$ too large to shut off)
4:  Find can’t start generators (those that have exceeded maximum allowed daily starts or are already ramping)
@@
6:  Order units in $\mathcal{G}_{d}$ into sets descending by $y_{i}$
7:  for each $\mathcal{G}_{d}$ that meets (Eq. 2b)-(Eq. 2e) do
8:   Solve economic dispatch Eq. 5 with $\mathcal{G}_{m},\mathcal{G}_{d}$ .
9:   if objective is lowest so far then
10:     Update best solution and objective.
11:   end if
12:  end for
13:  save Best solution and best objective.
14:end for
@@
Results from Algorithm 1 are not be directly comparable to fully coupled approach on computational time. Since we have shown our algorithm scales sub-quadratically [27], we believe it is possible to proceed with this approach and expand it in the future to include a fully time-coupled objective. In this work, the runtime constraints are still time-coupled and the system is forced to converge while fully accounting for minimum runtime or the number of starts.
Unlike [27], we do not compare RRUC results to MIP results. The large system sizes, combined with new constraints, make MIP prohibitively computationally intensive for the available hardware (Apple M3 Pro, 18GB RAM). Instead we focus specifically on RRUC and adopt a new metric to assess the efficiency of the algorithm: percentage increase in the objective function and percentage increase in runtime as a function of increase in the system size. If the objective function and runtime do not explode as the system size increases, the algorithm maintains its scaling properties and is efficient.
@@ III-B Simulation and Results
We multiply this generator set by 22 to pass 192 GW, using a random seed to rescale each generator in order to avoid the symmetry effects [19]. The noisy copies of the original 42 generators have a shift in capacity parameters of $\pm 10\%$ and a shift in daily starts or runtime hour parameters of $\pm 1$. For various system sizes, we further use multiples of this system.
@@
We present the results of the simulation in Fig. 2 using the same software and hardware as before. Since the reference MIP solver is too slow to be used as a benchmark for RRUC, in this study we are mainly interested in analyzing the scaling and performance of RRUC. The runtime and the objective function increase slightly faster than linearly with the number of generators. Specifically, the runtime increases on average about $2.6\times$ for every doubling of the system size. The $\approx O(n^{1.37})$ shows that the resulting systems are less complicated than those arising from a typical optimization problem which have $O(n^{1.5})$ scaling. For a general optimization problem, $O(n^{1.5})$ scaling is believed to be ideal, due to the linear solver bottleneck from the  [20].
@@ IV-A Mathematical Formulation
We now extend the problem to larger and more varied thermal power systems, which have have non-thermal units with intermittent characteristics. To operate such systems, we must account for system behavior at a higher resolution, which requires a sub hour granularity of unit commitment. Ramping is another important aspect of generator behavior at high temporal resolutions.
@@
To address the ramp constraints in the model, we use a different set of variables. Instead of the original $u_{i}$ that indicated that a generator is on or off, we allow the generator to be off, preparing (for ramp up), ramping up, on, or ramping down. For the model purposes, we assume that a generator can only change its status in one direction. If a generator is ramping up or preparing, it is assigned a variable $v_{i}=1$. After the required ramp time in each stage, its status will switch to on. If a generator is ramping down, it will eventually switch off and enter the set of generators that can start.
@@
$+\sum_{i\in G_{d}}\left[\left(a_{i}P_{i}^{2}+b_{i}P_{i}+c_{i}\right)u_{i}+K_{i}(1-u_{i}+v_{i})^{2}\right]$
@@
  • $u_{i}\in\{0,1\}$ for $i\in\mathcal{G}_{d}$: commitment status of discretionary generator $i$. $u_{i}=0$ is off and $u_{i}=1$ is on.
  • $v_{i}\in\{0,1\}$ for $i\in\mathcal{G}_{d}$ when $u_{i}$ was previously 1. $v_{i}=0$ represents staying off and $v_{i}=1$ is starting to ramp up.
@@
  • $P_{\max,i},P_{\min,i}$: maximum capacity and minimum stable output, determined by a unit’s physical characteristics.
@@
  • $K_{i}$: commitment change penalty for generator $i$.
  • $r_{d,i}$ and $r_{u,i}$: generator $i$’s limits for ramp down normalized so that $\Delta t=1$.
  • $\beta$: bias factor for selecting more efficient generators to meet the contingencies. In our model we set it to $0.001$.
  • $P_{typ,i}$, a typical production level of a generator during a peak event, which we define as $(4P_{\max,i}+P_{\min,i})/5$.
@@
  • $S_{r}$ - supply produced by generators that are already ramping before the solve (they cannot be changed) [^3] For generators ramping up, they are assumed to generate nothing for some number of periods, and then increase linearly to $P_{\min,i}$ in some number of time periods specific to the generator. Generators ramping down are assumed to decrease linearly from $P_{\min,i}$ in some number of time periods specific to the generator..
  • $S_{\max,r}$ total capacity of ramping up units (ramping down generators will be off before the contingency)
  • $S_{\min,r}$ minimum working capacity of ramping up units
@@
3:  Increment ramping periods for ramping units
4:  Switch unit that finished preparation to ramping up
5:  Turn units that have finished ramping up to on
6:  Turn units that have finished ramping down to off
7:  Find must run units (with minimum on time not yet reached or with $P_{t-1,i}-P_{min,i}$ too large to shut off)
8:  Find can’t start generators (those that have exceeded maximum allowed daily starts or are already ramping)
9:  Solve Eq. 6 with $u_{i},v_{i}\in\{0,1\}\rightarrow y_{i},z_{i}\in[0,1]$
10:  Order units in $\mathcal{G}_{d}$ into sets descending by $y_{i}$, then $z_{i}$
11:  for Each $\mathcal{G}_{d}$ satisfying (Eq. 6b)-(Eq. 6f) do
12:   Solve economic dispatch Eq. 5 with $\mathcal{G}_{m},\mathcal{G}_{d}$
13:   if Objective is lowest so far then
14:     Update best solution and objective
15:   end if
16:  end for
17:  save Best solution and best objective
18:end for
@@ IV-B Simulation and Results
Much of the data used in this Section has already been presented in Section III for runtime constraints. The additional assumptions about equipment characteristics are presented in Table II. While ramp characteristics usually come in MW/min, we use the % of $P_{\max,i}$ per minute as the original input into the simulation. We compile the percentage values the from the academic reports of Deutsches Institut für Wirtschaftsforschung [3] and the manufacturer catalog of GE Vernova [13]. The provided numbers, although summarising different sources, may still present the upper bound of ramp rates, as comprehensive survey of large thermal units suggest that in real systems power plants do not always deliver the nominal ramp rates specified for equipment [25].
@@
If the system starts with all generators off, Eq. 6 is usually infeasible. We initialized the system with the fastest ramping generators that can supply the previous demand. This may be suboptimal, but it is not the focus of this paper. Operators can initialize the system based on their own cold start protocol to a state that is known to be stable. If operators are warm starting using Algorithm 2 they can use the current system state.
@@
Fig. 4 summarizes the runtime and objective results. These results are comparable to those found in Section III using the same software and hardware. The runtime scaled roughly as $O(n^{1.6})$, where $n$ is the number of generators. Ideal scaling is believed to be bounded below by $O(n^{1.5})$, due to the scaling of the underlying linear system solvers [20]. This suggests that our optimization solver has close to ideal scaling. It is well below the exponential cost of MIP solvers. The average cost of the system did not increase with the system size. So adding extra generators does not result in scale cost effects.
@@ V Unit Commitment With Smooth Ramp Up
Figure 5: Unit $i$’s typical production cycle in Section V. When on, $P_{i}\in[P_{\min,i},P_{\max,i}]$, though it cannot exceed its ramping power. It can turn off if $P_{t-1,i}-r_{d,i}\leq P_{\min,i}$. The only other decision can be made when the unit is off, to start ramping up. During other periods, it is uncontrollable.
@@ VI Conclusion and Further Work
We proposed an algorithm to solve unit commitment with mechanical and ramping constraints by relaxing the integer constraints and then rounding solutions to satisfy all constraints. Our algorithm shows $\approx O(n^{1.6})$ time scaling in the number of generators. As the problem size increases, it retains accuracy nearly perfectly.